2. Consider the following data on the company Thor.

SPS = Sales per share 5.70

EPS = Earnings per share 2.75

DPS = Dividends per share 1.25 (this is last year dividend or D0 )

BV = Book value per share 5.10

NPM = Net profit margin 7.2%

ROE = Return on equity 18.5%

MP = Market price per share 26.29

Normal growth in EPS of 3.5%.

Super growth rate of 8.5% for 12 years.

2A. Calculate Current Price to Sales Ratio (multiple).

Explain what it means.

Calculate Current Price to Book Value Ratio (multiple).

Explain what it means.

Calculate the Current P/E

Explain what it means.

Calculate the current P/S

Explain what it means.
Answer #1
Ratio Name Ratio Formula Values Ratio
Price to Sales Ratio Market price per share / Sales per share =26.29/5.7 4.61
Price to Sales Ratio is a valuation metrics which measures the share price in times of sales per share, it means how much you pay price per share compared to sales per share.
Price to Book Ratio Market price per share / Book value per share =26.29/5.1 5.15
Price to Book Ratio compares the book value per share to price per share, It gives idea that how many times of book value share is trading in market
P/E Market price per share / Earnings per share =26.29/2.75 9.56
Price to earning (P/E) ratio is calculated to know for each dollar earned by company how many dollars investors are ready to pay. Basically it's a relationship between market share price and earnings per share
P/S Market price per share / Sales per share =26.29/5.7 4.61
Price to Sales Ratio is a valuation metrics which measures the share price in times of sales per share, it means how much you pay price per share compared to sales per share.